if the kinetic energy of an electron is given as k joules with an uncertainty of 6*10 ^*-6 then calculate the expression of uncertainty in velocity ,position . express your answer in terms of k only

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
Uncertainty (mistake) in KE(mv^2/2) = 6*10^-6 
So mistake in speed = 3*10^-6 
Instability primary... 
dP.dx
dP (blunder in force) = dV = 3*10^-6 
Figure dX 
Dx = (h/4π)/3x 10^-6 
(Estimations of h and π are consistent where h is board's steady and π=3.14) 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)